The demand functions for the only two goods in the economy are Q1 = 10 - 2p1 + p2 and Q2 = 10 - 2p2 + p1. There are five units of each good. Solve for the equilibrium: p1, p2, Q1, and Q2?
1
Expert's answer
2021-10-12T10:08:34-0400

Q1 = 10 - 2p+ p2

Q2 = 10 - 2p+ p1

There are five units of each good that means:

Q1 = 5

Q2 = 5

Put value of Qin first equation:

5 = 10 - 2p+ p2

2p- p2 = 5 ----------------(i)

Put value of Qin second equation:

5 = 10 - 2p+ p1

2p- p1 = 5 ----------------(ii)

then

p= 2p2 - 5

Put value of pin equation (i)

2(2p- 5) - p2 = 5

4p2 - 10 - p2 = 5

3p2 = 15

p2 = 15 / 3 = 5

p= 2*5 - 5 = 5

So, the equilibrium prices and quantities are:

Q1 = 5

Q2 = 5

p1 = 5

p2 = 5